When two given functions, 𝑓(π‘₯) = π‘₯ 2 + 3 π‘Žπ‘›π‘‘ 𝑔(π‘₯) = 2π‘₯ βˆ’ 5, are divided with each other, a new third function is obtained. Find (𝑓/𝑔)(π‘₯) and identify its domain


1
Expert's answer
2021-09-27T23:54:47-0400
"(f/g)(x)=\dfrac{x^2+3}{2x-5}"

"2x-5\not=0=>x\not=\dfrac{5}{2}"

"(f/g)(x)=\dfrac{x^2+3}{2x-5}"


"Domain: (-\infin, \dfrac{5}{2})\cup(\dfrac{5}{2}, \infin)"